OpenStudy (brittany):

John has a total of 13 coins in his pocket. Of the 13 coins, they are all either quarters or dimes. The total value of the coins is $2.50. How many dimes does John have in his pocket?

OpenStudy (curry):

5 dimes

OpenStudy (anonymous):

put x = number of quarters and therefore 13-x is number of dimes solve \[25x+10(13-x)=250\] for x

OpenStudy (anonymous):

\[25x+130-10x=250\] \[15x+130=250\] \[15x=120\] \[x=8\] so 8 quarters. the rest are dimes

OpenStudy (anonymous):

If the following equation is used, then no follow on subtraction is required. \[(13-x)(.25) +(x)(.10)=2.50\]